Determine three consecutive odd integers if the square of the largest integer is 33 less than the sum of the squares of the two smaller integers.
Answer by CubeyThePenguin(3113)

consecutive odd integers: (x-2), x, (x+2)

(x+2)^2 = (x-2)^2 + x^2 + 33
x^2 + 4x + 4 = x^2 - 4x + 4 + x^2 + 33
x^2 + 4x + 4 = 2x^2 - 4x + 37
0 = x^2 - 8x + 33

This equation has no integer solutions.